OpenStudy (anonymous):

Please help? What is the solution set of |x – 2| + 4 = 3?

jimthompson5910 (jim_thompson5910):

subtract 4 from both sides to get |x – 2| + 4 = 3 |x – 2| + 4-4 = 3-4 |x – 2| = -1 do you see what to do from here?

jimthompson5910 (jim_thompson5910):

keep in mind that |x| is either 0 or some positive number since the absolute value represents distance so |x-2| is either 0 or some positive number

OpenStudy (anonymous):

I don't know what to do next? Do I add 2 to both sides?

OpenStudy (jamesj):

Well, since |x-2| must be zero or positive, for what values of x is it the case that |x-2| = -1 ?

jimthompson5910 (jim_thompson5910):

|x-2| = -1 is basically saying "x is some number on the number line that is -1 units away from 2" is that possible?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\(|x-2|=-1\) stipulates \(x-2=-1\) and \(2-x=-1\).

jimthompson5910 (jim_thompson5910):

since it's not possible, there are no solutions

OpenStudy (anonymous):

Whoops, that's right, no possible solutions. Durr.

OpenStudy (anonymous):

Oh! I see now. No wonder it didn't make sense... Ha ha thanks, guys!